OpenStudy (anonymous):

what is the reciprocal of [(1/c) - (1/d)]?

OpenStudy (anonymous):

It is not c- d

OpenStudy (anonymous):

\[\frac{ 1 }{ c }-\frac{ 1 }{ d } = \frac{ d }{ cd } - \frac{ c }{ cd }\] \[=\frac{ d-c }{ cd }\] So the reciprocal should be: \[\frac{ cd }{ d-c }\]